write the equation of the parabola in standard form. find the coordinates of points corresponding to P and Q.
*i do not really know how to explain this question but here is my work if it helps. my teacher marked it wrong and i do not know what i did wrong.
f(x)= -3x^2 - 1x - 1
P' = (-2, 3)
Q'= (-1, 0)
*my teacher circled my answer for Q. i'm confused.

Answer by stanbon(75887)

write the equation of the parabola in standard form. find the coordinates of points corresponding to P and Q.
*i do not really know how to explain this question but here is my work if it helps. my teacher marked it wrong and i do not know what i did wrong.
f(x)= -3x^2 - 1x - 1
P' = (-2, 3)
Q'= (-1, 0)
*my teacher circled my answer for Q. i'm confused.
==================================
If you equation is correct.
f(-2) = -3(-2)^2-(-2) -1 = -12 + 2 -1 = -11
giving you (-2,-11) as a point on the parabola.
------------------------------------
Also, f(-1) = -3(1) -(-1) -1 = -3+1-1 = -3
giving you (-1,-3)
----------------------------
Cheers,
Stan H.